OpenStudy (anonymous):

Which steps could be used to solve this story problem? Krissy had 32 bagels. She gave an equal number of bagels to each child. There were 5 boys and 3 girls. How many bagels did the girls get altogether? A. Multiply 32 by 5. Multiply 32 by 3. Add the two products together. B. Add 5 and 3. Then multiply the sum by 32. C. Subtract 3 from 5. Then divide 32 by the difference. D. Add 5 and 3. Divide 32 by this sum. Then multiply the quotient by the number of girls.

OpenStudy (anonymous):

do anyone know this please will fan and medal!!!

OpenStudy (anonymous):

@LegendarySadist can u help me

OpenStudy (anonymous):

So if you want to find out how much she gave to the girls, you first have to find out how much she gave to everybody. Since she had 32 and she gave an equal amount to all 8, you'd have the equation \[\frac{ 32}{ 8 }=x\] Now just divide to get the amount taht everybody got.

OpenStudy (anonymous):

i got 20

OpenStudy (anonymous):

For x?

OpenStudy (anonymous):

i dont know do u know

OpenStudy (anonymous):

Well what is 32 divided by 8?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

is it B

OpenStudy (anonymous):

So every student gets 4. So if there are three girls, you just do \[4*3=?\] for the answer. Now that we've solved through it, which letter fits the way we did it?

OpenStudy (anonymous):

IS IT b

OpenStudy (anonymous):

B is saying to do 8 * 32. Did we ever do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Then it isn't our answer.

OpenStudy (anonymous):

then is it D I Really dont know then

OpenStudy (anonymous):

Yes, it is D. \[\frac{ 8 }{ 32 }*3\]

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

back u there

OpenStudy (anonymous):

im thinking D because you need to divide!